# Non-Hermitian Linear Electro-Optic Effect Through Interactions of Free and Bound Charges

### v1

@@ Abstract
In recent years, there has been significant interest in non-Hermitian phenomena, particularly optical gain, stemming from electro-optic effects. Typically, these effects are attributed to nonlinear interactions arising from the “anomalous velocity” of Bloch electrons. Here, we introduce a novel mechanism for achieving optical gain through nonlinear interactions between free and bound electrons. We propose a semiclassical phenomenological microscopic model that encompasses a broad spectrum of nonlinear interactions, consistent with passivity and with the principles of microscopic reversibility. Our analysis shows that under non-equilibrium conditions and within low-symmetry systems, the linearized electromagnetic response can display both nonreciprocal and gain properties. We examine the stability of these electrically-biased systems and illustrate their potential applications, including the development of optical isolators and traveling wave amplifiers.
@@ I Introduction
Remarkably, the optical gain arising from the non-Hermitian electro-optic (NHEO) effect may exhibit chiral properties, typically governed by the polarization handedness. Specifically, the interactions between the medium and the wave are polarization-dependent: one particular polarization causes the material to act as a regular dissipative system, while an orthogonal polarization state induces gain Lannebère et al. (2022); Rappoport et al. (2023); Morgado et al. (2024); Lannebère et al. (2025a). Previous studies have shown that such indefinte-gain responses can be leveraged to develop innovative photonic devices with unique characteristics, including optical isolators with gain, chiral lasers, and polarization-dependent mirrors Lannebère et al. (2022); Morgado et al. (2024); Lannebère et al. (2025b); Hakimi et al. (2024).
@@
As previously mentioned, the study of the NHEO effect has been centered on the nonlinear interactions involving drifting (free) electrons based on the anomalous velocity. In this paper, we expand on this by demonstrating that the NHEO effect can also emerge from the nonlinear coupling between bound and free electrons. Our analysis employs a semiclassical model, wherein free electrons are governed by standard transport and continuity equations, and bound charges adhere to Lorentzian dynamics. These dynamics are nonlinearly coupled through local fields, with the constraint that the interaction must satisfy to the principles of microscopic reversibility and passivity. We illustrate that under a static electric bias, the system’s linearized response is inherently non-Hermitian and nonreciprocal, with the type of response controlled by the symmetry of the material. This analysis paves new avenues for engineering robust NHEO effects in realistic materials.
@@
The article is organized as follows. In section II, we introduce a phenomenological nonlinear model for a material with free electrons possessing an anomalous velocity component. We demonstrate that the linearized optical response provided by the model agrees qualitatively with the results derived from Boltzmann theory. In section III, the model is extended to include both free and bound electrons. We analyze the possible forms of nonlinear coupling consistent with passivity and microscopic reversibility. We derive the general expression for the linearized permittivity response of the material. In section IV, we discuss various spatial symmetries compatible with nonreciprocal and gain responses, focusing on spatial symmetries of the 2mm group. We show that the NHEO effect in such systems originates a robust nonreciprocal and non-Hermitian electromagnetic response tailored by the electric field bias, even in the absence of an anomalous velocity term. In Section V, we explore wave propagation and phenomena associated with the proposed models. We show that materials of the 2mm symmetry group may enable the design of optical isolators and optical amplifiers. Conclusions are drawn in Sec. VI.
@@ II Phenomenological model for the NHEO effect with Berry curvature dipoles
where $q=-e$ is the electron charge, $m$ is an effective mass, $\Gamma$ is a collision frequency, and $\bf{p}$ may be understood as a crystal quasi-momentum. It is supposed that due to the interactions with the ionic lattice the electron velocity and the crystal quasi-momentum are related as follows:
@@
where $\overline{\zeta}$ is some system dependent tensor. The structure of the equations (1) and (2) is reminiscent of the semiclassical equations for Bloch electrons in a crystal with a nontrivial Berry curvature Xiao et al. (2010). In particular, the nonlinear interaction of the electrons with the crystal lattice originates a contribution to the velocity which is proportional to the electric field $\bf{E}$. In the context of Bloch electrons, this term is known as the anomalous velocity contribution and depends on the geometry (Berry curvature) of the crystal bands Xiao et al. (2010).
@@
×$\nabla\times{\bf{E}}=-{\mu_{0}}{\partial_{t}}{\bf{H}},\quad\quad\nabla\times{\bf{H}}={\varepsilon_{\rm{b}}}{\partial_{t}}{\bf{E}}{\rm{+}}{\bf{j}}_{\rm{c}}$ (3)
@@
${\bf{j}}_{\rm{c}}={n_{0}}q{\bf{v}}$ (4)
where $n_{0}$ is the electron density and $\varepsilon_{\rm{b}}$ is the (non-dispersive) permittivity of the background due to the bound electrons.
@@ II.1 Passivity and conservation laws
$\frac{d}{{dt}}\left[{\frac{{{n_{0}}}}{{2m}}{\bf{p}}\cdot{\bf{p}}}\right]+\frac{{{n_{0}}\Gamma}}{m}{\bf{p}}\cdot{\bf{p}}={\bf{E}}\cdot{\bf{j}}_{\rm{c}}.$ (5)
From the Poynting theorem Kong (1986), it is well-known that $\nabla\cdot{\bf{S}}+{\partial_{t}}{W_{{\rm{EM}}}}=-{\bf{E}}\cdot{\bf{j}}_{\rm{c}}$, with ${\bf{S}}={\bf{E}}\times{\bf{H}}$ the Poynting vector and ${W_{{\rm{EM}}}}=\frac{1}{2}{\varepsilon_{\rm{b}}}{\bf{E}}\cdot{\bf{E}}+\frac{1}{2}{\mu_{\rm{0}}}{\bf{H}}\cdot{\bf{H}}$ the electromagnetic energy density. Combining this result with Eq. (5), it is found that the proposed phenomenological model adheres to the conservation law:
$\nabla\cdot{\bf{S}}+{\partial_{t}}{W_{{\rm{tot}}}}=-Q,\quad\quad Q=\frac{{{n_{0}}\Gamma}}{m}{\bf{p}}\cdot{\bf{p}},$ (6)
${W_{{\rm{tot}}}}=\frac{1}{2}{\varepsilon_{\rm{b}}}{\bf{E}}\cdot{\bf{E}}+\frac{1}{2}{\mu_{\rm{0}}}{\bf{H}}\cdot{\bf{H}}+{n_{0}}\frac{1}{{2m}}{\bf{p}}\cdot{\bf{p}}.$ (7)
@@ II.2 Linearized dynamical response
Let us now consider the scenario where the material is biased by a static electric field ${\bf{E}}_{0}$ created by some external generator. We write the total field as the sum the DC component and a weak dynamic field: ${\bf{E}}={{\bf{E}}_{0}}+\left({{{\bf{E}}_{\omega}}{e^{-i\omega t}}+\rm{c.c.}}\right)$. As Eq. (1) is linear, the corresponding quasi-momentum has a similar decomposition, ${\bf{p}}={{\bf{p}}_{0}}+\left({{{\bf{p}}_{\omega}}{e^{-i\omega t}}+\rm{c.c.}}\right)$ with
@@
Combining Eqs. (2) and (4), it is evident that the (linearized) dynamical current is given by ${{\bf{j}}_{{\rm{c,}}\omega}}=\frac{{{n_{0}}q}}{m}\left[{{{\bf{p}}_{\omega}}+\left({\overline{\zeta}\cdot{{\bf{p}}_{\omega}}}\right)\times{{\bf{E}}_{0}}+\left({\overline{\zeta}\cdot{{\bf{p}}_{0}}}\right)\times{{\bf{E}}_{\omega}}}\right]$. Then, from Eq. (8), it follows that the linearized current is related to the dynamical field as ${{\bf{j}}_{{\rm{c,}}\omega}}=\left({{{\overline{\sigma}}_{\rm{D}}}+{{\overline{\sigma}}_{{\rm{EO}}}}}\right)\cdot{{\bf{E}}_{\omega}}$, where the ${{{\overline{\sigma}}_{\rm{D}}}+{{\overline{\sigma}}_{{\rm{EO}}}}}$ is the optical conductivity. It consists of a Drude-like response (${{\overline{\sigma}}_{\rm{D}}}$) and an electro-optic response that originates from the static-bias (${{\overline{\sigma}}_{\rm{EO}}}$):
${\overline{\sigma}_{\rm{D}}}={\varepsilon_{0}}\omega_{p}^{2}\frac{1}{{\Gamma-i\omega}}{{\bf{1}}_{3\times 3}},\quad\quad\quad{\overline{\sigma}_{{\rm{EO}}}}={\varepsilon_{0}}\omega_{p}^{2}\left[{\frac{1}{\Gamma}\left({\overline{\zeta}\cdot{{\bf{E}}_{0}}}\right)\times{{\bf{1}}_{3\times 3}}+\frac{{-1}}{{\Gamma-i\omega}}{{\bf{E}}_{0}}\times\overline{\zeta}}\right]$ (9)
@@
Interestingly, the term ${{\overline{\sigma}}_{\rm{EO}}}$ shares the same structure as the electro-optic conductivity derived using the Boltzmann formalism in Morgado et al. (2024). Indeed, the results from the phenomenological model align precisely with those calculated using Boltzmann’s approach when the tensor $\overline{\zeta}$ satisfies:
@@
Here, $\overline{\mathbf{D}}$ represents the Berry curvature dipole of the material. This similarity confirms that the phenomenological model accurately describes the NHEO effect. As extensively discussed in prior research Morgado et al. (2024); Rappoport et al. (2023), an electrically biased material may exhibit optical gain. The phenomenological model elucidates how this property is consistent with the material’s global passivity [Eq. (6)] and highlights that the optical gain originates from the nonlinear interactions between DC and dynamic fields.
@@ II.3 Lagrangian formulation
Importantly, as demonstrated in Appendix A, linearizing any system described by a Lagrangian subjected to a time-independent bias—essentially, a system in equilibrium—inevitably results in a Hermitian linearized response. Consequently, the system of equations (1)-(4) cannot be derived from a Lagrangian density, even when approaching the limit $\Gamma\to 0^{+}$. More broadly, we conclude that a nonlinear conservative system in equilibrium cannot exhibit gain upon linearization, thereby precluding its use for implementing the NHEO effect. Consistent with this finding, the analyses in Morgado et al. (2024); Rappoport et al. (2023) demonstrate that the Berry curvature dipole can only be nonzero in metallic systems—specifically, systems where a static electric bias induces a drifting current, representing a nonequilibrium scenario.
@@
In this context, it is noteworthy that the anomalous velocity contribution in Eq. (2) is represented by a tensor lacking transposition symmetry, $-\frac{1}{m}{\bf{E}}\times\overline{\zeta}$. In contrast, the term describing the effective mass is always a symmetric tensor (as exemplified here by $\frac{1}{m}{{\bf{1}}_{3\times 3}}$, which corresponds to a scalar). Thus, the correction due to the anomalous velocity cannot be regarded as a field-dependent effective mass. This underscores the unique characteristics of the (nonlinear) term associated with the NHEO effect.
@@ III Model of free and bound charges nonlinearly coupled
Next, we explore alternative physical mechanisms that may potentially contribute to the NHEO effect, beyond the influence of the anomalous velocity. Specifically, we demonstrate that the NHEO effect can emerge from the interaction between nonlinearly coupled free and bound charges, even without the presence of an anomalous velocity term. Our microscopic model adheres to the principles of microscopic reversibility and passivity, even though the system dynamics is not described by a Lagrangian.

### v2

@@ Abstract
In recent years, there has been growing interest in non-Hermitian phenomena in low-symmetry conductors, particularly optical gain driven by electro-optic effects. Conventional semiclassical treatments typically attribute these effects to nonlinear interactions associated with the anomalous velocity of Bloch electrons. Here, we present a phenomenological microscopic model that not only recovers these anomalous-velocity contributions, but also incorporates interband effects that become significant at higher frequencies. Our model captures a wide range of nonlinear interactions while remaining consistent with passivity and microscopic reversibility. Using this broader framework, we study the nonlinear interactions between free and bound electrons as an alternative mechanism for optical gain. We show that, under non-equilibrium conditions in low-symmetry conductors, the linearized electromagnetic response can exhibit both nonreciprocity and gain, even without anomalous velocity contributions. Finally, we analyze the stability of electrically biased systems and highlight potential applications such as optical isolators and traveling-wave amplifiers.
@@ I Introduction
Moreover, metallic or semimetallic systems without inversion symmetry are known to exhibit giant nonlinear optical responses Wu et al. (2017); Patankar et al. (2018); Hu et al. (2023). For example, Ref. Wu et al. (2017) reports that transition metal monopnictides such as TaAs have a $\mathbf{\overline{\chi}}^{(2)}$ nonlinearity nearly an order of magnitude larger than that of GaAs, a widely used material in nonlinear optics. This enhancement originates from the itinerant nature of the free carriers, which enables significantly larger displacements under optical excitation than in dielectrics, where electrons are strongly bound to the atomic nuclei. A potential drawback, however, is that metallic-like systems may exhibit higher dissipation than dielectric materials.
@@
Remarkably, the optical gain arising from the non-Hermitian electro-optic (NHEO) effect may exhibit chiral properties, typically governed by the polarization handedness. Specifically, the interactions between the medium and the wave are polarization-dependent: one particular polarization causes the material to act as a regular dissipative system, while an orthogonal polarization state induces gain Lannebère et al. (2022); Rappoport et al. (2023); Morgado et al. (2024); Lannebère et al. (2025a). Previous studies have shown that such indefinite-gain responses can be leveraged to develop innovative photonic devices with unique characteristics, including optical isolators with gain, chiral lasers, and polarization-dependent mirrors Lannebère et al. (2022); Morgado et al. (2024); Lannebère et al. (2025b); Hakimi et al. (2024).
@@
While the anomalous velocity contribution to electro-optic effects in low-symmetry conductors such as Weyl semimetals is already well understood, a phenomenological framework for capturing other contributions—particularly those involving interactions between free and bound electrons—remains underdeveloped. In this work, we introduce a minimal Lorentz-type phenomenological model that describes such interactions. Heuristically, the proposed couplings are expected to capture nonlinear effects arising from interband processes in low-symmetry conductors. Our analysis employs a semiclassical model, wherein free electrons are governed by standard transport and continuity equations, and bound charges follow Lorentzian dynamics. These dynamics are nonlinearly coupled through local fields, with the constraint that the interaction must satisfy to the principles of microscopic reversibility and passivity. We illustrate that under a static electric bias, the system’s linearized response is inherently non-Hermitian and nonreciprocal, with the type of response controlled by the symmetry of the material. Our model provides a foundation for interpreting nonlinear responses in low-symmetry conductors and for fitting experimental data as it becomes available.
@@
The article is organized as follows. In section II, we introduce a phenomenological nonlinear model for a material with free electrons possessing an anomalous velocity component. We demonstrate that the linearized optical response provided by the model agrees qualitatively with the results derived from semiclassical Boltzmann theory. In section III, the model is extended to include both free and bound electrons. We analyze the possible forms of nonlinear coupling consistent with passivity and microscopic reversibility. We derive the general expression for the linearized permittivity response of the material. In section IV, we discuss various spatial symmetries compatible with nonreciprocal and gain responses, focusing on spatial symmetries of the 2mm point group and on the electro-optic response arising from interband interactions. We show that the NHEO effect in such systems originates a robust nonreciprocal and non-Hermitian electromagnetic response tailored by the electric field bias, even in the absence of an anomalous velocity term. In Section V, we explore wave propagation and phenomena associated with the proposed models. We show that materials belonging to the 2mm point group may enable the design of optical isolators and optical amplifiers. Conclusions are drawn in Sec. VI.
@@ II Phenomenological model for the NHEO effect with Berry curvature dipoles
where q=eq=-e is the electron charge, $m$ is an effective mass, $\Gamma$ is a collision frequency, and $\bf{p}$ may be understood as a crystal quasi-momentum. It is supposed that due to the interactions with the ionic lattice the electron velocity and the crystal quasi-momentum are related as follows:
@@
where $\overline{\zeta}$ is some system dependent tensor. The structure of the equations (1) and (2) is reminiscent of the semiclassical equations for Bloch electrons in a crystal with a nontrivial Berry curvature Xiao et al. (2010). In particular, the nonlinear interaction of the electrons with the crystal lattice originates a contribution to the velocity which is proportional to the electric field $\bf{E}$. In the context of Bloch electrons, this term is known as the anomalous velocity contribution and depends on the geometry (Berry curvature) of the crystal bands Xiao et al. (2010). Equation (2) is intended to effectively model the contribution of the Bloch electrons near the Fermi level, and thereby its structure differs slightly from that of the conventional anomalous velocity.
@@
$\nabla\times{\bf{E}}=-{\mu_{0}}{\partial_{t}}{\bf{H}},\quad\quad\nabla\times{\bf{H}}={\varepsilon_{\text{b}}}{\partial_{t}}{\bf{E}}{\text{ + }}{\bf{j}}_{\text{c}}$ (3)
@@
${\bf{j}}_{\text{c}}={n_{0}}q{\bf{v}}$ (4)
where $n_{0}$ is the electron density and $\varepsilon_{\text{b}}$ is the (non-dispersive) permittivity of the background due to the bound electrons.
@@ II.1 Passivity and conservation laws
$\frac{d}{{dt}}\left[{\frac{{{n_{0}}}}{{2m}}{\bf{p}}\cdot{\bf{p}}}\right]+\frac{{{n_{0}}\Gamma}}{m}{\bf{p}}\cdot{\bf{p}}={\bf{E}}\cdot{\bf{j}}_{\text{c}}.$ (5)
From the Poynting theorem Kong (1986), it is well-known that $\nabla\cdot{\bf{S}}+{\partial_{t}}{W_{{\text{EM}}}}=-{\bf{E}}\cdot{\bf{j}}_{\text{c}}$, with ${\bf{S}}={\bf{E}}\times{\bf{H}}$ the Poynting vector and ${W_{{\text{EM}}}}=\frac{1}{2}{\varepsilon_{\text{b}}}{\bf{E}}\cdot{\bf{E}}+\frac{1}{2}{\mu_{\text{0}}}{\bf{H}}\cdot{\bf{H}}$ the electromagnetic energy density. Combining this result with Eq. (5), it is found that the proposed phenomenological model adheres to the conservation law:
$\nabla\cdot{\bf{S}}+{\partial_{t}}{W_{{\text{tot}}}}=-Q,\quad\quad Q=\frac{{{n_{0}}\Gamma}}{m}{\bf{p}}\cdot{\bf{p}},$ (6)
${W_{{\text{tot}}}}=\frac{1}{2}{\varepsilon_{\text{b}}}{\bf{E}}\cdot{\bf{E}}+\frac{1}{2}{\mu_{\text{0}}}{\bf{H}}\cdot{\bf{H}}+{n_{0}}\frac{1}{{2m}}{\bf{p}}\cdot{\bf{p}}.$ (7)
@@ II.2 Linearized dynamical response
Let us now consider the scenario where the material is biased by a static electric field ${\bf{E}}_{0}$ created by some external generator. We write the total field as the sum the DC component and a weak dynamic field: ${\bf{E}}={{\bf{E}}_{0}}+\left({{{\bf{E}}_{\omega}}{e^{-i\omega t}}+\text{ c.c.}}\right)$. As Eq. (1) is linear, the corresponding quasi-momentum has a similar decomposition, ${\bf{p}}={{\bf{p}}_{0}}+\left({{{\bf{p}}_{\omega}}{e^{-i\omega t}}+\text{c.c.}}\right)$ with
@@
Combining Eqs. (2) and (4), it is evident that the (linearized) dynamical current is given by ${{\bf{j}}_{{\text{c,}}\omega}}=\frac{{{n_{0}}q}}{m}\left[{{{\bf{p}}_{\omega}}+\left({\overline{\zeta}\cdot{{\bf{p}}_{\omega}}}\right)\times{{\bf{E}}_{0}}+\left({\overline{\zeta}\cdot{{\bf{p}}_{0}}}\right)\times{{\bf{E}}_{\omega}}}\right]$. Then, from Eq. (8), it follows that the linearized current is related to the dynamical field as ${{\bf{j}}_{{\text{c,}}\omega}}=\left({{{\overline{\sigma}}_{\text{D}}}+{{\overline{\sigma}}_{{\text{EO}}}}}\right)\cdot{{\bf{E}}_{\omega}}$, where the ${{\overline{\sigma}}_{\text{D}}}+{{\overline{\sigma}}_{{\text{EO}}}}$} is the optical conductivity. It consists of a Drude-like response (${\overline{\sigma}}_{\text{D}}$}) and an electro-optic response that originates from the static-bias (${\overline{\sigma}}_{\text{EO}}$}):
${\overline{\sigma}_{\text{D}}}={\varepsilon_{0}}\omega_{p}^{2}\frac{1}{{\Gamma-i\omega}}{{\bf{1}}_{3\times 3}},\quad\quad\quad{\overline{\sigma}_{{\text{EO}}}}={\varepsilon_{0}}\omega_{p}^{2}\left[{\frac{1}{\Gamma}\left({\overline{\zeta}\cdot{{\bf{E}}_{0}}}\right)\times{{\bf{1}}_{3\times 3}}+\frac{{-1}}{{\Gamma-i\omega}}{{\bf{E}}_{0}}\times\overline{\zeta}}\right]$ (9)
@@
Interestingly, the term ${\overline{\sigma}}_{\text{EO}}$} shares the same structure as the electro-optic conductivity derived using the semiclassical Boltzmann formalism in Morgado et al. (2024). Indeed, the results from the phenomenological model align precisely with those calculated using Boltzmann’s approach when the tensor $\overline{\zeta}$ satisfies:
@@
Here, $\overline{\mathbf{D}}$ represents the Berry curvature dipole of the material. This similarity confirms that the phenomenological model describes precisely the electro-optic response predicted by the semiclassical Boltzmann theory. As extensively discussed in prior research Morgado et al. (2024); Rappoport et al. (2023), the electrically biased material may exhibit optical gain. Our phenomenological model elucidates how this property is consistent with the material’s global passivity [Eq. (6)] and highlights that the optical gain originates from the nonlinear interactions between DC and dynamic fields.
@@ II.3 Lagrangian formulation
Importantly, as demonstrated in Appendix A, linearizing any system described by a Lagrangian subjected to a time-independent bias—essentially, a system in equilibrium—inevitably results in a Hermitian linearized response. Consequently, the system of equations (1)-(4) cannot be derived from a Lagrangian density, even when approaching the limit $\Gamma\to 0^{+}$. More broadly, we conclude that a nonlinear conservative system in equilibrium cannot exhibit gain upon linearization, thereby precluding its use for implementing the NHEO effect. Consistent with this finding, the analyses in Morgado et al. (2024); Rappoport et al. (2023) demonstrate that the Berry curvature dipole can only be nonzero in metallic systems, specifically, systems where a static electric bias induces a drifting current, representing a nonequilibrium scenario.
@@
In this context, it is noteworthy that the anomalous velocity contribution in Eq. (2) is represented by a tensor lacking transposition symmetry, $\frac{1}{m}{\bf{E}}\times\overline{\zeta}$. In contrast, the term describing the effective mass is always a symmetric tensor (as exemplified here by $\frac{1}{m}{{\bf{1}}_{3\times 3}}$, which corresponds to a scalar). Thus, the correction due to the anomalous velocity cannot be regarded as a field-dependent effective mass. This underscores the unique characteristics of the (nonlinear) term associated with the NHEO effect.
@@ III Model of free and bound charges nonlinearly coupled
Next, we explore alternative physical mechanisms that may potentially contribute to the NHEO effect, beyond the influence of the anomalous velocity. Specifically, we demonstrate that the NHEO effect can emerge from the interaction between nonlinearly coupled free and bound charges, even without the presence of an anomalous velocity term. The motivation for considering such interactions is to model interband effects of the electro-optic susceptibility response $\mathbf{\overline{\chi}}_{\text{EO}}$. Indeed, the $\mathbf{\overline{\chi}}_{\text{EO}}$ computed from first principle approaches, such as via the Kubo formalism Aversa and Sipe (1995); Jia et al. (2024), can be decomposed as $\mathbf{\overline{\chi}}_{\text{f-f}}(\omega)+\mathbf{\overline{\chi}}_{\text{f-b}}(\omega)$ , where $\mathbf{\overline{\chi}}_{\text{f-f}}(\omega)=\mathbf{\overline{\sigma}}_{\text{EO}}(\omega)/(-i\omega\varepsilon_{0})$ represents the piece of the electro-optic response associated with the Berry curvature dipole (semiclassical response, already discussed in Sect. II.2), whereas $\mathbf{\overline{\chi}}_{\text{f-b}}(\omega)=\mathbf{\overline{\sigma}}_{\text{EO}}^{\text{\,inter}}(\omega)/(-i\omega\varepsilon_{0})$ represents the remaining part of the electro-optic response due to interband interactions. Note that for low frequencies, the semiclassical part of the response dominates because $\lim_{\omega\to 0^{+}}\mathbf{\overline{\sigma}}_{\text{EO}}^{\text{\,inter}}(\omega)=0$. However, for high frequencies, the additional interband contribution can become relevant, particularly near resonances. Our objective here is to introduce a phenomenological model that can describe $\mathbf{\overline{\chi}}_{\text{f-b}}(\omega)$.
Our microscopic model adheres to the principles of microscopic reversibility and passivity, even though the system dynamics is not described by a Lagrangian. As further discussed later, in principle, the parameters of the phenomenological model can be mapped to the (electro-optic) second-order susceptibilities obtained from band theory Jia et al. (2024). The magnitude of such interband effects in low-symmetry conductors Wu et al. (2017); Patankar et al. (2018); Hu et al. (2023) may, in principle, be comparable to those of intraband NHEO effects (and, perhaps, even supplant them near resonances), making their study relevant for applications.